A rocket rises vertically with acceleration of 3.2m/s^2. It runs out of fuel at 750 m. It’s acceleration from here is gravity.
1- what is the velocity of the rocket when it runs out of fuel.
2- how long does it take to reach 750 m.
3- what is the highest altitude the rocket reaches.
4- what velocity does it strike earth.
5- how long is it in the air
1
Expert's answer
2020-06-08T10:14:06-0400

Explanations & Calculations


  • The rocket starts from rest & after 750 m it enters a vertical motion under the gravity.
  • Therefore, until 750 m rocket is under 3.2 acceleration & just after 750 m its under the 9.8 ms-2 gravity.


1) Velocity after 750 m: apply V2 = U2 +2as vertically upward

"\qquad\qquad
\begin{aligned}
\small V^2 & = \small 0^2 + 2\times 3.2ms^{-2} \times 750 m\\
\small V & = \small \sqrt{4800m^2s^{-2}}\\
V &= \small \bold{40\sqrt{3} =69.28ms^{-1}} 
\end{aligned}"


2) Time taken to reach the 750 m : apply V = U +at

"\qquad\qquad
\begin{aligned}
\small 40\sqrt{3}ms^{-1} & = \small 0 +3.2\times t\\
\small t &= \small \bold{\frac{25\sqrt{3}}{2}=21.65s}
\end{aligned}"


3) Maximum height it reaches from 750 m : apply V = U +2as vertically upward (a = g =9.8 ms-2 )

"\qquad\qquad
\begin{aligned}
\small 0^2 &= \small (40\sqrt3ms^{-1})^2 + 2\times(-9.8ms^{-2}) \times h\\
\small h &= \small \frac{4800m^2s^{-2}}{19.6ms^{-2}}\\
&= \small 244.898m
\end{aligned}"

Therefore, the highest altitude it reaches = 750 m +244.898 m = 994.898 m


4) Velocity, it strikes the ground with : apply V2 = U2 + 2as vertically downward from the 750 m height.

"\qquad\qquad
\begin{aligned}
\small V^2 & = \small (-40\sqrt3 )^2 + 2\times (+9.8)\times(+750 m)\\
\small V &= \small \sqrt{19500 m^2s^{-2}}\\
&= \small \bold{10\sqrt{195}= 139.64 ms^{-1}}
\end{aligned}"


5) Time spent in free fall ,t : apply S = Ut +0.5st2 vertically downward from the height 750 m.

"\qquad\qquad
\begin{aligned}
\small +750 m &= \small (-40\sqrt3)\times t +\frac{1}{2} \times9.8 \times t^2\\
\small 0 &= \small 9.8t^2-80\sqrt3t-1500\\
\small t &= \small 21.32 s \,(\text{or} -7.18s )\\

\end{aligned}"

Therefore, total time airborne T = 21.65s + 21.32s = 42.97s